A particle follows a spiral orbit given by r = c (theta)2 under an unknown force law. Prove that such an orbit is possible in a central field. Also find the force law?

1
Expert's answer
2021-03-25T14:07:56-0400

Answer

u="\frac{1}{r}=\frac{1}{c\theta^2}"


For finding force law

"\frac{d^2u}{d\theta^2}+u=-\frac{Fm}{l^2u^2}"

Putting value of

"\frac{d^2u}{d\theta^2}=\frac{6}{c\theta^4}"

And u in above equation

"\frac{6}{c\theta^4}+\frac{1}{c\theta^2}=-\frac{Fm}{l^2u^2}"

Changing in r form

Then force law become

"F=-\frac{m}{l^2}(6c+r)"